# GEMS: Group Emotion profiling through Multimodal Situational understanding

### v1

@@ 1 Introduction
[…]
@@ 2 Background
Fig. 2: Data statistics for the GEMS-VGAF dataset. The figure presents the class-wise distribution of individual emotions, group emotions, and situational contexts in the proposed dataset. Note, Quar: Quarrelling, CasGat: Casual Gathering, GrpAct: Group Activities and CelPar:Celebration Party. The y-axis indicates the number of samples corresponding to each class.
@@ 3 VGAF-GEMS
Situational Emotion. For situation-based emotion labeling, we categorized each video clip into one of ten distinct scene types following literature [5]: Quarrelling, Meeting, Funeral, Sports, Casual-gathering, Group-activities, Protest, Celebration-party, Fighting, and Show. According to social sciences [30], we use the multi-lable classification of different situational emotion such as Active, Enthusiastic, Anxious, Angry, Attentive, Distressed, Sad, Relaxed, Rejected and Happy. Please note that it is manually annotated. We have also conducted a label space survey with 10 participants for annotating situational emotion and situation. The average perception level agreement for 10 videos are 69% for situation and 67% for situational emotion, respectively.
@@
VGAF-
GEMS
Zero-shot Video-LLM [28] - - - - - 01.17 0.0105 0 0.0075 0 15.77 12.70
@@
Dataset Statistics Class-wise Distribution of label space We split the dataset into three nonoverlapping sets: train and validation (similar to Original VGAF partition). In each set, we have the following statistics: 2,661 training, 766 validation videos (5 sec duration). The overview of classwise relabelled data distribution is presented in Figure 2.
@@
Prompt 3.1: Video-ChatGPT Interaction System: Initialize Video-ChatGPT. Human: {<<<Interaction>>> What are the interactions between the people in the video? Your answer should be one or multiple of the following: interactions categories. Please think and list all possible answers.} AI: {EXAMPLE OUTPUT is the interaction class e.g. insults, thanks, helps, disagrees with, walks with, jokes, gives, compliments etc.} Human: {<<<Scene>>> What is the background of the groups of people in the video? Your answer should be one of the following: Scenes categories. Please think and generate only one word as the answer.} AI: {EXAMPLE OUTPUT is scene class.} Human: {<<<Event>>> What are the event of the group of people in the video? Your answer should be one or multiple of the following: event categories} AI: {EXAMPLE OUTPUT is event class.} Human: {<<<Relationship>>> What are the relationship of the group of people in the video? Your answer should be one or multiple of the following: relationship categories} AI: {EXAMPLE OUTPUT is relationship class.}
@@ 4 Benchmark and Metrics
Evaluation Matrices. Following the standard evaluation protocol [37, 1, 17, 4], we utilize class-wise accuracy (for discrete emotion), CCC and MSE (for continuous emotion) as evaluation matrices. Please note that VGAF test set labels are not publicly available. Thus, we utilize only train and validation partition.
@@ 5 Results and Discussion
[…]

### v2

@@ 1 Introduction
  • We perform comprehensive analysis and benchmark the proposed dataset utilizing state-of-the-art emotion recognition algorithms. We benchmark the datasets for two different learning paradigms, i.e., zero-shot and fully supervised. The significant drop in performance (i.e., $\sim\textbf{54.80}$ w.r.t VGAF dataset with supervised framework) directly indicate that our dataset will be a valuable asset to pursue further research in this domain.
@@ 2 Background
Fig. 2: Data statistics for the VGAF-GEMS dataset. The figure presents the class-wise distribution of individual emotions, group emotions, and situational contexts in the proposed dataset. Note, Quar: Quarrelling, CasGat: Casual Gathering, GrpAct: Group Activities and CelPar:Celebration Party. The y-axis indicates the number of samples corresponding to each class.
@@ 3 VGAF-GEMS
Situational Emotion. For situation-based emotion labeling, we frame this as a multi-label classification task. Following literature [5], each video clip is categorized into one of ten distinct situations: Quarrelling, Meeting, Funeral, Sports, Casual-gathering, Group-activities, Protest, Celebration-party, Fighting, and Show. Based on social sciences [30], we use a corresponding set of ten situational emotions: Active, Enthusiastic, Anxious, Angry, Attentive, Distressed, Sad, Relaxed, Rejected, and Happy. To create the ground truth for training, each situation is mapped to its predefined set of expected emotions, which is then encoded as a 10-dimensional multi-hot vector. For a video labeled with a specific situation, the positions in the vector for its mapped emotions were set to 1, while all others were set to 0. For instance, a video categorized as ’Quarrelling’ receives a target vector where the indices for ’Active’, ’Enthusiastic’, ’Anxious’, and ’Angry’ are set to 1. The situation labels were manually annotated, and a label space survey with 10 participants showed an average perception agreement of 69% for the situation and 67% for the situational emotion, respectively.
@@
VGAF- GEMS Zero-shot Video-LLM [28] - - - - - 01.17 0.0105 0 0.0075 0 15.77 12.70
@@
Dataset Statistics Class-wise Distribution of label space We split the dataset into two nonoverlapping sets: train and validation (similar to Original VGAF partition). In each set, we have the following statistics: 2,661 training, 766 validation videos (5 sec duration). The overview of classwise relabelled data distribution is presented in Figure 2.
@@
Prompt 3.1: Video-ChatGPT Interaction System: Initialize Video-ChatGPT. Human: {<<Interaction>> What are the interactions between the people in the video? Your answer should be one or multiple of the following: interactions categories. Please think and list all possible answers.} AI: {EXAMPLE OUTPUT is the interaction class e.g. insults, thanks, helps, disagrees with, walks with, jokes, gives, compliments etc.} Human: {<<Scene>> What is the background of the groups of people in the video? Your answer should be one of the following: Scenes categories. Please think and generate only one word as the answer.} AI: {EXAMPLE OUTPUT is scene class.} Human: {<<Event>> What are the event of the group of people in the video? Your answer should be one or multiple of the following: event categories} AI: {EXAMPLE OUTPUT is event class.} Human: {<<Relationship>> What are the relationship of the group of people in the video? Your answer should be one or multiple of the following: relationship categories} AI: {EXAMPLE OUTPUT is relationship class.}
@@ 4 Benchmark and Metrics
Evaluation Metrics. Following the standard evaluation protocol [37, 1, 17, 4], we utilize class-wise accuracy (for discrete emotion), CCC and MSE (for continuous emotion) as evaluation metrics. Please note that VGAF test set labels are not publicly available. Thus, we utilize only train and validation partition.
@@ 5 Results and Discussion
Quantitative Analysis. The results of proposed VGAF-GEMS benchmark is presented in Table 1. The zero shot capability of the state-of-the-art video-LLM is quite weak. For group level inference, it performs around 1.17% for 9-class group emotion, 15.77% for situational emotion understanding. In supervised benchmark as well, the model performs notably low in multi task learning. Please note that the addition of S3 attention module boosts the performance quite a lot ($9.46\%\rightarrow 54.80\%$). Please note that we are the first to propose the emotional compositionality over video in terms of individual, group and situational emotion. Thus, publicly available benchmarks are not comparable.